OpenStudy (anonymous):

The measure of one angle of a triangle is 1degree less than twice the measure of the second angle, and the measure of the third angle is 21degrees more than twice the sum of the measure of the other two. Find the measure of the third angle

OpenStudy (anonymous):

a = 2b - 1 c = 2(a + b) + 21 a + b + c = 180

OpenStudy (anonymous):

Substitute the right side of the first equation into "a" in the 2nd and 3rd equations. Work with the altered equations 2 and 3 as 2 equations in 2 unknowns of "b" and "c".